Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 60mph and train B is traveling at 80mph. Train A passes a station at 10:25 pm, if train B passes the same station at 10:40 pm, at what time will train B catch up to train A?
Answer by mananth(16946)

Train A 60 kmph
Train B 80 kmph
Train A 10:25
Train B 10:40
Difference in time= 00:15 0.25
Train A will have covered 15 miles beforeTrain B starts from station
catch up distance= 15 miles
catch up speed = 80 -60 mph
catch up speed = 20 mph
Catchup time = catchup distance/catch up speed
catch up time= 15 / 20
catch up time= 0.75 hours
catch up time= 45 minutes
10:40 + 0:45= 11:25 pm